OpenStudy (anonymous):

Solve x2 - 7x = -13.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x^2 - 7x = -13 }\) \(\large\color{black}{ \displaystyle x^2 - 7x +(7/2)^2= -13 +(7/2)^2}\) and so forth, ... (completing the square)

OpenStudy (anonymous):

A. \[x=\frac{ -7\pm i \sqrt{6} }{ 2 }\] B. \[x=\frac{ -7\pm i \sqrt{3} }{ 2 }\] C. \[x=\frac{ 7\pm i \sqrt{6} }{ 2 }\] D. \[x=\frac{ 7\pm i \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

@SolomonZelman